thomaster (thomaster):

Three replicate determinations of the signal for a standard solution of an analyte at a concentration of 10.0 ppm give values of 0.163, 0.157, and 0.161 (arbitrary units), respectively. The signal for a method blank was found to be 0.002. Calculate the concentration of analyte in a sample that gives a signal of 0.118. So confused :S Can somebody explain this?

OpenStudy (missmob):

this is analytical chemistry right?

thomaster (thomaster):

yes

OpenStudy (missmob):

its hard

thomaster (thomaster):

If it was easy, I wouldn't ask for help :)

thomaster (thomaster):

I think i know how to calculate it.. but what to do with the different signal values. average them and subtract the blank? That way it would be: AVG(0.163,0.157,0,161) = 0.160 subtract the blank = 0.160-0.002 = 0.158 10 ppm = 0.158 ?? ppm = 0.118 \[\frac{10*0.118}{0.158}~=7.47~ ppm\]

thomaster (thomaster):

oh forgot to subtract the blank from the sample signal: \[\frac{10*0.116}{0.158}~=7.34~ ppm\] Can somebody confirm that i used the correct method?

thomaster (thomaster):

I ruled out the error from the average signal of the standard solution, not the signal of the sample. The correct answer is 7.33 so i wasn't that far away from it :)

OpenStudy (chmvijay):

yaaa i think you are right :) i just gone through it now :)

thomaster (thomaster):

great :) tnx